# Uncertainty Modeling in Graph Neural Networks via Stochastic Differential Equations

### v3

@@ A.4 GNSDE as a Continuous Representation of Graph ResNet with Stochastic Noise Insertion
Here, $\Delta t$ represents a fixed time step and $\Delta\mathbf{W}_{j}$ is a Brownian increment, normally distributed with mean zero and variance $\Delta t$. This numerical discretization is analogous to a Graph Residual Network (Graph ResNet) with a specific structure, where Brownian noise is injected at each residual layer. Therefore, the Graph Neural SDE can be interpreted as a deep Graph ResNet where the depth corresponds to the number of discretization steps of the SDE solver.
@@ Appendix B Details of The Experimental Setup
Table 2: Uniform Hyperparameters of the LGNSDE, GNODE, Bayesian GCN, GCN, Ensemble of GCN models for all datasets.
@@
Method SRK SK4 n/a
Dropout 0.2 0.2 0.2
Diffusion g 1.0 n/a n/a
Metric Model Cora Citeseer Computers Photo Pubmed
MICRO-AUROC (↑) GCN 0.9654 ± 0.0050 0.9173 ± 0.0068 0.9680 ± 0.0016 0.9905 ± 0.0003 0.9006 ± 0.0139
GNODE nan ± nan 0.9146 ± 0.0063 0.9569 ± 0.0067 0.9885 ± 0.0007 0.8857 ± 0.0203
LGNSDE (Our) 0.9667 ± 0.0036 0.9111 ± 0.0072 0.9691 ± 0.0032 0.9909 ± 0.0004 0.9007 ± 0.0091
AURC (↓) GCN 0.9966 ± 0.0007 0.9966 ± 0.0011 0.9994 ± 0.0005 0.9987 ± 0.0015 0.9994 ± 0.0004
GNODE nan ± nan 0.9967 ± 0.0011 0.9994 ± 0.0004 0.9998 ± 0.0001 0.9915 ± 0.0163
BGCN 0.9972 ± 0.0004 0.9963 ± 0.0010 0.9994 ± 0.0002 0.9989 ± 0.0005 0.9996 ± 0.0004
ENSEMBLE 0.9970 ± 0.0012 0.9967 ± 0.0012 0.9994 ± 0.0002 0.9989 ± 0.0006 0.9996 ± 0.0005
LGNSDE (Our) 0.9970 ± 0.0003 0.9971 ± 0.0005 0.9995 ± 0.0003 0.9997 ± 0.0002 0.9995 ± 0.0005
Accuracy (↑) GCN 0.8105 ± 0.0173 0.7258 ± 0.0137 0.8098 ± 0.0048 0.9116 ± 0.0021 0.7570 ± 0.0229
GNODE nan ± nan 0.7235 ± 0.0159 0.7911 ± 0.0098 0.9053 ± 0.0032 0.7577 ± 0.0231
LGNSDE (Our) 0.8113 ± 0.0128 0.7120 ± 0.0119 0.8247 ± 0.0103 0.9169 ± 0.0021 0.7595 ± 0.0168
Table 3: AUROC (Mean ± Std), AURC (Mean ± Std), and Accuracy (Mean ± Std) for all datasets. Red denotes the best-performing model, and blue denotes the second-best-performing model.

### v4

@@ A.4 LGNSDE as a Continuous Representation of Graph ResNet with Stochastic Noise Insertion
Here, $\Delta t$ represents a fixed time step and $\Delta\mathbf{W}_{j}$ is a Brownian increment, normally distributed with mean zero and variance $\Delta t$. This numerical discretization is analogous to a Graph Recurrent Network (Graph ReNet) with a specific structure, where Brownian noise is injected at each recurrent layer. Therefore, the Graph Neural SDE can be interpreted as a deep Graph ReNet where the depth corresponds to the number of discretization steps of the SDE solver.
@@ Appendix B Details of The Experimental Setup
B.1 Hyperparameter Search
Table 6: Hyperparameter Grid Search Configuration
Hyperparameter Values
Learning Rate {0.001, 0.005, 0.01, 0.1}
Weight Decay {0.01, 0.001, 0.0005, 0.0001}
Epoch {15, 100, 200, 300}
Dropout {0.0, 0.1, 0.3, 0.5}
Hidden Dimension {16, 32, 64, 128, 256}
Step Size {0.01, 0.05, 0.1, 0.2}
Table 7: Hyperparameters left out of the grid search for all models and used for all datasets.
@@ B.1 Hyperparameter Search
Method SRK RK4 N/A
Early Stop 20 20 20
Diffusion $\mathbf{G}_{\mathcal{G}}$ 1.0 N/A N/A
B.2 Active Learning
Table 8: Dataset statistics before and after active learning.
Dataset # Nodes # Links Training/Validation/Test Split Initial # Training Labels Final # Training Labels
Cora 2,708 5,429 140/500/1000 140 280
Citeseer 3,327 4,732 120/500/1000 120 240
Computers 13,752 245,861 200/500/1000 200 400
Photo 7,650 119,081 160/500/1000 160 320
Pubmed 19,717 44,338 60/500/1000 60 120
B.3 LGNSDE Hyperparameters
Table 9: Accuracy vs. prior drift in Cora.
Prior drift Accuracy (%)
-100.0 81.29
-10.0 82.21
-5.0 80.52
-1.0 82.17
-0.5 83.66
0.0 82.74
0.5 82.58
1.0 82.41
5.0 80.48
10.0 80.80
20.0 82.74
100.0 82.17
Table 10: Accuracy vs. prior diffusion in Cora.
Sigma ($\sigma$) Accuracy (%)
0.1 79.45
0.2 80.13
0.5 81.32
0.8 80.52
1.0 80.79
1.5 78.93
2.0 78.47
5.0 76.85
7.0 68.78
10.0 60.54
Appendix C More Experiments and Benchmarks
C.1 LGNSDE backbone model GCN vs GAT
Metric Model Cora Citeseer
AUROC (↑) LGNSDE-GAT 0.8249 ± 0.0321 0.8003 ± 0.0488
LGNSDE-GCN 0.7614 ± 0.0804 0.8258 ± 0.0418
AURC (↓) LGNSDE-GAT 0.0108 ± 0.0011 0.0536 ± 0.0080
LGNSDE-GCN 0.0168 ± 0.0070 0.0479 ± 0.0109
Accuracy (↑) LGNSDE-GAT 0.9474 ± 0.0002 0.8642 ± 0.0090
LGNSDE-GCN 0.9471 ± 0.0003 0.8729 ± 0.0108
Figure 4: Illustrating two Drift function backbone the standard GCN vs Graph Attention Network.
C.2 Computational Cost Comparison
Model Time per Epoch (s)
GCN $0.19$
GAP $0.23$
BGCN $0.36$
GNODE $0.38$
GRAPH GP $0.72$
GNSD $0.92$
ENSEMBLE $1.13$
LGNSDE (Ours) 1.23
Table 12: Time Taken per Epoch for Different Models .
C.3 Further Benchmarking and Complexities
Table 13: OOD detection performance comparison on Cora with different OOD constructions.
Model Label leave-out Feature perturbation
AUROC AUPR in AUPR out FPR95 DET ACC AUROC AUPR in AUPR out FPR95 DET ACC
GCN 88.67 74.52 95.87 52.40 78.80 75.40 54.22 86.69 80.98 69.94
GAT 90.81 71.97 96.50 44.91 83.65 89.39 80.19 92.80 70.37 76.16
GRAND 88.58 75.72 94.28 63.59 82.34 87.04 74.81 93.42 55.35 76.43
GREAD 88.40 73.80 95.01 57.35 80.74 86.49 70.59 93.01 57.10 75.30
MSP 91.58 80.59 96.72 44.03 84.51 90.58 82.31 92.02 50.33 80.45
ODIN 49.24 24.27 75.45 100.00 49.95 49.80 26.92 72.95 100.00 49.98
Mahalanobis 66.83 36.69 85.92 82.96 59.62 60.46 40.65 74.52 99.59 58.52
GNNsafe 92.68 82.03 97.48 31.54 83.48 93.28 88.16 96.35 43.43 83.71
GCN-Ensemble 90.97 80.35 97.37 29.92 85.73 89.01 79.71 94.57 61.74 74.57
BGCN 91.16 79.41 96.60 46.30 84.40 84.82 75.40 91.20 75.20 78.09
GKDE 82.80 65.12 92.62 70.59 71.64 80.22 63.05 90.01 75.01 64.74
GPN 90.10 79.98 96.13 50.71 82.81 91.89 82.62 96.01 42.32 81.56
GNSD 94.76 88.45 97.73 27.38 89.74 91.77 94.41 88.23 26.27 86.92
LGNSDE (Ours) 93.14 85.01 96.65 49.03 81.42 90.61 86.35 92.77 31.63 83.05
Table 14: OOD detection comparison on Amazon-Computers with different OOD constructions.
Model Label leave-out Feature perturbation
AUROC AUPR in AUPR out FPR95 DET ACC AUROC AUPR in AUPR out FPR95 DET ACC
GCN 82.35 56.46 93.67 56.06 74.72 80.55 78.53 78.55 80.67 75.09
GAT 80.66 53.19 93.05 53.91 72.65 73.69 78.00 65.61 97.41 75.76
GRAND 80.27 52.51 92.84 54.81 71.99 84.93 81.29 87.33 54.98 65.34
GREAD 80.56 54.05 92.70 54.14 72.68 85.38 79.07 87.60 59.10 68.29
MSP 74.88 47.53 89.64 75.52 68.85 72.86 74.50 67.73 95.70 70.81
ODIN 71.78 37.70 89.87 70.54 50.18 79.13 80.09 77.09 83.09 66.75
Mahalanobis 71.87 37.76 89.87 70.24 50.18 74.47 67.54 76.28 82.48 50.04
GNNsafe 90.50 77.20 95.05 48.25 84.47 89.46 95.17 84.01 75.62 76.49
GCN-Ensemble 79.53 52.39 91.99 69.28 73.51 77.71 79.45 72.60 94.20 77.51
BGCN 82.19 57.52 93.30 57.43 73.55 83.60 82.93 81.50 72.49 75.78
GKDE 76.46 48.18 90.64 73.36 64.35 71.69 71.40 69.04 90.83 69.70
GPN 88.76 68.23 96.45 42.08 81.02 87.92 85.99 85.98 67.10 81.24
GNSD 94.06 82.27 97.06 31.47 88.76 95.95 94.76 94.62 15.69 91.28
LGNSDE (Ours) 90.71 74.25 97.78 68.20 82.28 90.86 93.64 84.00 35.21 82.49
Table 15: Time and Memory Complexity of Models
Model Time Complexity Memory Complexity
Graph Posterior Network (GPN) $O(N\cdot K)$ $O(N\cdot C)$
Graph Gaussian Processes (GGP) $O(NM^{2})$ $O(NM)$
Graph Neural ODEs (GNODE) $O(E\cdot F^{2}\cdot\text{NFEs})$ $O(1)$
Latent Graph Neural SDEs (LGNSDE) $O(L\log L(|E|d+|V|d))$ $O(1)$
Table 16: Explanation of Time and Memory Complexity Components
Component Description
$N$ Number of nodes in the graph
$K$ Average degree of the nodes (used in GPN)
$M$ Number of inducing points (used in GGP to reduce complexity)
$E$ Number of edges in the graph
$F$ Feature dimensionality for nodes or edges
NFEs Number of function evaluations during ODE solving (for GNODE)
$L$ Number of steps required by the SDE or ODE solver (for LGNSDE)
$C$ Number of classes in classification tasks (for GPN)